# Can GPT models Follow Human Summarization Guidelines? A Study for Targeted Communication Goals

### v2

@@ Abstract
This study investigates the ability of GPT models (ChatGPT, GPT-4 and GPT-4o) to generate dialogue summaries that adhere to human guidelines. Our evaluation involved experimenting with various prompts to guide the models in complying with guidelines on two datasets: DialogSum (English social conversations) and DECODA (French call center interactions). Human evaluation, based on summarization guidelines, served as the primary assessment method, complemented by extensive quantitative and qualitative analyses. Our findings reveal a preference for GPT-generated summaries over those from task-specific pre-trained models and reference summaries, highlighting GPT models’ ability to follow human guidelines despite occasionally producing longer outputs and exhibiting divergent lexical and structural alignment with references. The discrepancy between ROUGE, BERTScore, and human evaluation underscores the need for more reliable automatic evaluation metrics.111The generated summaries and human annotations will be made publicly available at https://github.com/yongxin2020/LLM-Sum-Guidelines.
@@ 1 Introduction
Although instruction-tuned Large Language Models (LLMs) have shown impressive performance in several benchmark datasets, they still struggle with various challenging tasks [1, 2, 3].
For instance, general-purpose models like ChatGPT generally underperform in summarization tasks compared to task-specific pre-trained models in automatic metrics [3, 4, 5]. In dialogue sumarization, ChatGPT demonstrates poorer performance than state-of-the-art (SOTA) fine-tuned models on the ROUGE metric for datasets like SAMSum [6] and DialogSum [7].
However, the underlying causes of GPT models’ underperformance on automatic summarization metrics remain poorly understood. Human summarization is a complex cognitive process that involves understanding, condensing, and conveying the essential information of a source text. During the data annotation phase, human approach summarization by adhering to the provided guidelines. Given that human annotators have to follow these guidelines when creating reference summaries, we speculate that the poor performance of LLMs could be attributed to the lack of explicit summarization instructions tailored to specific communication goals. This is particularly critical for dialogue summarization, where targeted communication goals play a significant role.
Even though dialogue summarization is a well-established task, it remains a challenging task due to the absence of a consensus for what constitutes an ideal dialogue summary [8]. The approach to dialogue summarization is heavily influenced by specific communication objectives, which vary across different contexts such as meetings, or customer service interactions. A review of guidelines from major dialogue summarization datasets reveals that each corpus defines its own set of objectives for creating reference summaries, applying different summary criteria to meet specific needs [9]. For instance, in the context of customer service, while TWEETSUMM [10] offers both extractive and abstractive summaries, CSDS [11] provides three distinct summaries for each dialogue: an overall summary and two role-oriented summaries (user and agent).
Our key contributions in evaluating GPT models’ adherence to human summarization guidelines are as follows:
@@
Evaluating three GPT models against task-specific pre-trained models on two multilingual datasets with distinct communication goals (DialogSum and DECODA);
@@ 2 Experimental Setup
DialogSum consists of social conversations in English and includes 12,460, 1,500 and 500 samples in its training, validation and test splits, respectively. It was recently used in a challenge [13], which favors comparison of results.
@@
We used OpenAI ChatGPT (gpt-3.5-turbo)222OpenAI GPT-3.5: https://platform.openai.com/docs/models/gpt-3-5, the latest version available used in September 2023 is gpt-3.5-turbo-0613, GPT-4333OpenAI GPT-4: https://platform.openai.com/docs/models/gpt-4, version gpt-4-0613, and GPT-4o444gpt-4o-2024-08-06: https://platform.openai.com/docs/models#gpt-4o APIs for our experiments.
gpt-3.5-turbo, the most capable GPT-3.5 variant, is optimized for conversational tasks while maintaining strong performance on standard completion tasks. The model handles inputs up to tokens, compared to GPT-4’s -token limit and GPT-4o’s -token context windows. To ensure a stable output, we configured the temperature parameter at while maintaining the default values for all other parameters.
For comparison with task-specific pre-trained models, we also fine-tuned BART-large555https://huggingface.co/facebook/bart-large on the DialogSum dataset and fine-tuned BARThez666https://huggingface.co/moussaKam/barthez on the DECODA dataset. Details of the training are presented in Appendix A.
@@
| Human Guideline (HG) for DialogSum | System: Write a summary based on following criteria: the summary should (1) convey the most salient information of the dialogue and; (2) be brief (no longer than 20% of the conversation length) and; (3) preserve important named entities within the conversation and; (4) be written from an observer perspective and; (5) be written in formal language. In addition, you should pay extra attention to the following aspects: 1) Tense Consistency: take the moment that the conversation occurs as the present time, and choose a proper tense to describe events before and after the ongoing conversation. 2) Discourse Relation: If summarized events hold important discourse relations, particularly causal relation, you should preserve the relations if they are also in the summary. 3) Emotion: you should explicitly describe important emotions related to events in the summary. 4) Intent Identification: Rather than merely summarizing the consequences of dialogues, you should also describe speakers’ intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation. User: [Test Dialogue] |
@@
The different prompts used in our experiments are given in Table 1. The WordLimit (WL) prompt simply constrains the word length of the output to be less than X words [1]. Another prompt we tested was to provide a text only version of the full Human Guideline (HG). For DialogSum, the annotation guideline was directly described in Chen et al. [7]. For DECODA, we contacted the authors [12] and obtained the exact guidelines used to write the synopses. In addition, to examine if giving the system a role helps, we proposed Human Guideline with Role (HGR), which begins with "You are an annotator …". We have also investigated a two-step iterative approach, called HG(R)WL, which first uses the summaries generated by HG or HGR as input, then re-injects them into the model to reduce the length of the output using the WL prompt.
@@
However, few metrics excel across all dimensions, leading recent studies to explore LLMs for evaluating text quality [18]. Following this trend, we adopted an LLMs-as-judge approach, using GPT-4o as the backbone for automatic evaluation on a subset of model predictions.
@@ 3.1 Quantitative Results
The results of the automatic evaluations on the test set are presented in Table 2. The performance of the ChatGPT model with the WordLimit prompt is approximately points lower for R1, points lower for R2, and points lower for RL compared to state-of-the-art pre-trained models. However, BERTScore suggests only a small semantic difference. The results for GPT-4 with the WordLimit prompt are slightly better than those of ChatGPT, except for R2. The results of GPT-4o are comparable to those of GPT-4, with the WL prompt demonstrating superior performance across most metrics, except for BERTScore.
Although the DialogSum guideline states "be brief (no longer than 20% of the conversation length)", the results of the HG prompt indicate that instructions tailored for human annotators may not be suitable when directly used as instructions to generate reference-like summaries with ChatGPT, as responses tend to be longer, leading to lower ROUGE and BERTScore values. This effect could have been reduced by specifying the role of the annotator in the prompt (HGR), which consistently improved performance across all measures.
@@
In terms of average length, the WordLimit prompt consistently gave the shortest summaries for both GPT models. GPT-4 and GPT-4o better followed the WordLimit prompt in obtaining a summary of less than words compared to ChatGPT.
@@
When comparing the GPT models, all three models performed better with the WordLimit prompt, while yielding similar results with the iterative prompt (HGRWL). In terms of average length, the WordLimit prompt consistently gave the shortest summaries for all GPT models. Again, the GPT-4 and GPT-4o models better followed the WordLimit prompt in obtaining a summary of less than words compared to ChatGPT, which always exceeded the limit.
@@
Using DeepSeek-R1999https://api-docs.deepseek.com/guides/reasoning_model as our evaluation backbone, we assessed the generated summaries against human summarization guideline criteria (see Appendix B for details and prompts). We evaluated a subset of 20 samples from the DECODA dataset, with results shown in Table 4.
However, we observed that even when the prompt explicitly provided guidelines for judging summaries and specified that only the note should be returned, many evaluation results contained not only scores but also scoring explanations – sometimes in French and sometimes in English.
The GPT models’ generated summaries achieved strong scores across all four criteria. The best results for Faithfulness, Main Issues, and Resolution came from 4-HGRWL, 3.5-WL, and 3.5-HGRWL respectively, while the References performed best on Sub-Issues. Considering average scores across all criteria, 4-HGRWL achieved the highest scores while Barthez obtained the lowest.
| Faithfulness (VN) | Main Issues (VN) | Sub-Issues (VN) | Resolution (VN) | Average (VN) | |
| Reference | 3.65 1.09 (20) | 4.35 1.09 (20) | 3.85 1.27 (20) | 3.15 1.81 (20) | 3.75 1.32 (20) |
| 3.5-WL | 4.00 0.86 (20) | 4.60 0.75 (20) | 3.60 1.10 (20) | 3.35 1.73 (20) | 3.89 1.11 (20) |
| 3.5-HGRWL | 3.20 1.47 (20) | 4.00 1.30 (20) | 3.75 1.21 (20) | 3.95 1.15 (20) | 3.73 1.28 (20) |
| 4-WL | 4.00 1.12 (20) | 4.40 0.82 (20) | 3.42 1.39 (19)† | 2.45 1.70 (20) | 3.57 1.26 (20) |
| 4-HGRWL | 4.15 0.93 (20) | 4.45 0.76 (20) | 3.80 0.83 (20) | 3.50 1.54 (20) | 3.98 1.02 (20) |
| BARThez | 1.70 1.13 (20) | 2.00 1.65 (20) | 1.42 0.84 (19)‡ | 1.60 1.27 (20) | 1.68 1.22 (20) |
@@ 3.3 Manual Error Analysis
We also analyzed the data points where GPT-generated summaries exhibit the greatest discrepancy, characterized by low ROUGE scores but high BERTScore values. For comparison, we selected predictions from two models: 4-WL, which uses a simple WordLimit prompt, and 4-HGRWL, which incorporates human summarization guidelines as an intermediate step. We identified the five examples with the largest discrepancies for DialogSum predicted by 4-WL (Table 6) and 4-HGRWL (Table 15). For DECODA, we present samples predicted by 4-WL (Table 7) and 4-HGRWL (Table 16), both translated from French into English.
@@
First, we look at the summaries generated by 4-WL. We observed in the DialogSum examples with the greatest discrepancies that the references specifically articulate by preserving mentions of #Person1# and #Person2#, summarizing the conversation while retaining named entities. In contrast, GPT-4 predictions often summarize the conversation at a higher, generalized level, using terms like "two friends" or "employees" instead of specific tags. This demonstrates that GPT-generated summaries fail to meet the following guideline: "3) Preserve important named entities within the conversation" when using a simplified prompt.
In addition, GPT-4 predictions don’t match the references concerning the guideline emphasizing "Intent Identification"1010104) Intent Identification: Rather than merely summarizing the consequences of dialogues, you should also describe speakers intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation. . For instance, in the second example, the prediction references only Person 2, omitting the intent of Person 1, who is interviewing the other individual. In the fourth example, neither Person 1 nor Person 2 is mentioned, and their intents are absent, as the prediction reduces the dialogue to a general remark about the weather. In the fifth example, the prediction fails to follow the "use person tags" rule, summarizing the conversation by identifying the interlocutors as "A Mexican businesswoman" and "her American-Colombian taxi driver". These variations probably influenced the quantitative results evaluated by automated metrics, as evidenced by the high BERTScore values indicating good semantic similarity, while the low ROUGE-L scores suggest minimal overlap in the longest common sequence. This disparity highlights the tendency of GPT-generated summaries using a simple length constraint prompt to capture the gist and key information of the conversation while diverging in lexical and structural alignment with reference summaries.
For the DECODA dataset, the discrepancy may stem from the distinct characteristics of the reference synopses. These summaries focus strictly on issues and resolutions, whereas GPT-generated predictions often adopt a more individual-centered perspective, using terms such as "un organisateur (an organizer)", "une cliente (a client)", and "un représentant (a representative)", etc. This difference results in relatively high BERTScore values, reflecting strong semantic similarity, but low ROUGE-L scores due to minimal word overlap with the references. Additionally, reference synopses are often extremely brief and sometimes incomplete sentences, as seen in the fourth example: "circulation du T2" (T2 circulation). Furthermore, all samples except the fourth ("T2 circulation") fail to conclude the resolution in the generated summaries.
@@
Next, we examine the generated summaries of 4-HGRWL. On the DialogSum dataset, we observe better retention of named entities and personal tags (e.g., Person 1) compared to 4-WL, indicating that the intermediate HGR step helps the model to follow specific rules from the human summarization guidelines. However, the issue of overlooking Intent Identification persists, as prediction often focus on one individual’s intent while omitting the other’s, whereas references typically indicate both interlocutors’ intents.
For DECODA, similar to the 4-WL outputs, most generated summaries employ an individual-centered perspective that creates divergence in -gram overlap and semantic similarity. While reference summaries maintain conciseness, model predictions tend toward detail while frequently omitting critical Resolution elements. For example, the first sample omits "Transfer to the relevant department" despite this being an explicit requirement in the summarization guidelines.
In summary, comparing the data points with the greatest discrepancy (low ROUGE-L but high BERTScore) between 4-WL and 4-HGRWL, the latter better adheres to guidelines regarding named entities and personal tags. However, both models share similar shortcomings: in DialogSum, they overlook "Intent Identification"; in DECODA, they use an individual-centered perspective and omit the "Resolution".

### v3

@@ Abstract
This study investigates the ability of GPT models (ChatGPT, GPT-4 and GPT-4o) to generate dialogue summaries that adhere to human guidelines. Our evaluation involved experimenting with various prompts to guide the models in complying with guidelines on two datasets: DialogSum (English social conversations) and DECODA (French call center interactions). Human evaluation, based on summarization guidelines, served as the primary assessment method, complemented by extensive quantitative and qualitative analyses. Our findings reveal a preference for GPT-generated summaries over those from task-specific pre-trained models and reference summaries, highlighting GPT models’ ability to follow human guidelines despite occasionally producing longer outputs and exhibiting divergent lexical and structural alignment with references. The discrepancy between ROUGE, BERTScore, and human evaluation underscores the need for more reliable automatic evaluation metrics.111The generated summaries and human annotations are publicly available at https://github.com/yongxin2020/LLM-Sum-Guidelines
@@ 1 Introduction
Although instruction-tuned Large Language Models (LLMs) have shown impressive performance on several benchmark datasets, they still struggle with various challenging tasks [1, 2]. On automatic metrics like ROUGE, general-purpose models (e.g., ChatGPT) consistently trail behind task-specific, fine-tuned models [3, 4, 5]. In dialogue summarization specifically, ChatGPT demonstrates poorer performance than state-of-the-art fine-tuned models on datasets including SAMSum [6] and DialogSum [7]. However, human evaluation studies reveal a more nuanced picture: instruction-tuned LLMs can demonstrate human-aligned summarization capabilities that automatic metrics may fail to capture [8].
The underlying causes of GPT models’ underperformance on automatic summarization metrics remain poorly understood. Human summarization is a complex cognitive process that involves understanding, condensing, and conveying essential information while adhering to specific guidelines. Given that human annotators have to follow these guidelines when creating reference summaries, we speculate that the performance discrepancy could be attributed to the lack of explicit summarization instructions tailored to specific communication goals during model prompting. This is particularly critical for dialogue summarization, where targeted communication goals play a significant role.
Even though dialogue summarization is a well-established task, it remains challenging due to the absence of a consensus for what constitutes an ideal dialogue summary [9]. The approach to dialogue summarization is heavily influenced by specific communication objectives, which vary across different contexts such as meetings, or customer service interactions. A review of guidelines from major dialogue summarization datasets reveals that each corpus defines its own set of objectives for creating reference summaries, applying different summary criteria to meet specific needs [10]. For instance, in the context of customer service, while TWEETSUMM [11] offers both extractive and abstractive summaries, CSDS [12] provides three distinct summaries for each dialogue: an overall summary and two role-oriented summaries (user and agent).
In this work, we evaluated the ability of GPT models to generate dialogue summaries that adhere to human guidelines, which provides the following contributions:
@@
Evaluating three GPT models against task-specific pre-trained models on English and French datasets with distinct communication goals (DialogSum and DECODA);
@@
2 Related Work
2.1 Evaluation of Summarization Tasks
The dominant approach for automatic summarization evaluation employs similarity-based metrics. ROUGE [13] remains the standard, measuring n-gram overlap (e.g., ROUGE-1, ROUGE-2, ROUGE-L) between system outputs and human references. To capture semantic similarity beyond lexical overlap, embedding-based metrics like BERTScore [14] leverage contextual embeddings. Alternative paradigms include natural language inference (NLI) for evaluating factual consistency through entailment [15] and question answering (QA)-based methods that use question generation and answering [16]. More recently, LLM-based evaluation has emerged, encompassing metrics derived from, prompted by, or fine-tuned on LLMs [17]. For instance, G-Eval [18] operates without references and assesses dimensions like coherence, consistency, fluency, and relevance.
Beyond automatic metrics, human evaluation of summaries has evolved into a multi-dimensional framework assessing faithfulness, coherence, relevance, completeness, and conciseness [19]. This framework is applied across diverse domains: from general news benchmarks evaluating faithfulness, coherence, and relevance [8], to specialized tasks like topic-focused dialogue summarization assessed for completeness, relevance, and factual consistency [20], to domain-specific evaluations (e.g., bookings, interviews) that employ key fact validation to calculate scores for faithfulness, completeness, and conciseness [21].
Nevertheless, a number of challenges remain. Automatic metrics show limited correlation with human judgment [19]. LLM-based approaches face challenges such as prompt sensitivity, limited confidence calibration, the need for human oversight (restricting full automation), and limited explainability [17]. Furthermore, the field lacks consensus on evaluation standards, with substantial variability in criteria and protocols across studies [22], complicating reliable and reproducible assessment.
2.2 Instruction Following in LLMs
Prompt engineering has emerged as an indispensable technique for guiding LLMs, enabling users to interact with generative AI systems through carefully designed instructions without modifying core model parameters [23, 24]. Systematic surveys have categorized numerous distinct prompt engineering techniques based on their targeted functionalities [23], though challenges persist including biases, factual inaccuracies, and interpretability gaps.
Research demonstrates that prompt formatting impacts LLM performance through various approaches. For example, the Chain-of-Thought (CoT) technique prompts models to generate intermediate reasoning steps before delivering a final answer [25]. Optimization by PROmpting (OPRO) employs an iterative process where the LLM generates new solutions from prompts containing previous solutions with their values, which are then evaluated and added to subsequent prompts [26]. Studies also show that emotional intelligence can be leveraged through EmotionPrompt, which combines original prompts with emotional stimuli to enhance performance [27]. Furthermore, the Rephrase and Respond approach allows LLMs to rephrase and expand human-posed questions before providing responses, serving as an effective method for improving output quality [28]. These techniques collectively demonstrate that prompt quality influences response quality, with even simple formatting adjustments proving effective for performance improvement.
@@ 3 Experimental Setup
DialogSum consists of social conversations in English and includes 12,460, 1,500 and 500 samples in its training, validation and test splits, respectively. It was used in a challenge [30], which favors comparison of results.
@@
We used OpenAI ChatGPT (gpt-3.5-turbo),222OpenAI GPT-3.5: https://platform.openai.com/docs/models/gpt-3-5, version gpt-3.5-turbo-0613 GPT-4,333OpenAI GPT-4: https://platform.openai.com/docs/models/gpt-4, version gpt-4-0613 and GPT-4o444gpt-4o-2024-08-06: https://platform.openai.com/docs/models#gpt-4o APIs for our experiments.
gpt-3.5-turbo is optimized for conversational tasks while maintaining strong performance on standard completion tasks. The model handles inputs up to tokens, compared to GPT-4’s -token limit and GPT-4o’s -token context windows. To ensure a stable output, we configured the temperature parameter at while maintaining the default values for all other parameters.
For comparison with task-specific pre-trained models, we also fine-tuned BART-large555https://huggingface.co/facebook/bart-large on the DialogSum dataset and BARThez666https://huggingface.co/moussaKam/barthez on the DECODA dataset. The details of the training are presented in Appendix A.
@@
| Human Guideline (HG) for DialogSum | System: Write a summary based on the following criteria: the summary should (1) convey the most salient information of the dialogue and; (2) be brief (no longer than 20% of the conversation length) and; (3) preserve important named entities within the conversation and; (4) be written from an observer perspective and; (5) be written in formal language. In addition, you should pay extra attention to the following aspects: 1) Tense Consistency: take the moment that the conversation occurs as the present time, and choose a proper tense to describe events before and after the ongoing conversation. 2) Discourse Relation: If summarized events hold important discourse relations, particularly causal relation, you should preserve the relations if they are also in the summary. 3) Emotion: you should explicitly describe important emotions related to events in the summary. 4) Intent Identification: Rather than merely summarizing the consequences of dialogues, you should also describe speakers’ intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation. User: [Test Dialogue] |
@@
The different prompts used in our experiments are given in Table 1. The WordLimit (WL) prompt simply constrains the word length of the output to be less than words [1]. Another prompt we tested was to provide a text only version of the full Human Guideline (HG). For DialogSum, the annotation guideline was directly described in Chen et al. [7]. For DECODA, we contacted the authors [29] and obtained the guidelines used to write the synopses. In addition, to examine if giving the system a role helps, we proposed Human Guideline with Role (HGR), which begins with "You are an annotator …". We have also investigated a two-step iterative approach, called HG(R)WL, which first uses the summaries generated by HG or HGR as input, then re-injects them into the model to reduce the length of the output using the WL prompt.
@@
However, few metrics excel across all dimensions, leading recent studies to explore LLMs for evaluating text quality [18]. Following this trend, we adopted an LLMs-as-judge approach, using DeepSeek-R1 as the backbone for automatic evaluation on a subset of model predictions.
@@ 4.1 Quantitative Results
Table 2 presents the automatic evaluation results on the test set. While the ChatGPT model with the WordLimit prompt underperforms state-of-the-art pre-trained models by approximately 15 (R1), 10 (R2), and 20 (RL) points, BERTScore indicates only a minor semantic discrepancy. The GPT-4 model with the same prompt shows a slight overall improvement over ChatGPT, except on R2. Furthermore, GPT-4o performs comparably to GPT-4, with the WL prompt yielding superior results on most metrics, though not on BERTScore.
Although the DialogSum guideline states "be brief (no longer than 20% of the conversation length)", the results of the HG prompt indicate that instructions tailored for human annotators may not be suitable when directly used as instructions to generate reference-like summaries with ChatGPT, as responses tend to be longer, leading to lower ROUGE and BERTScore values. This effect could have been reduced by specifying the role of the annotator in the prompt (HGR), which consistently improved performance with reference to HG across all measures.
@@
In terms of average length, GPT-4 and GPT-4o better followed the WordLimit prompt in obtaining a summary whose length is closest to the reference compared to ChatGPT.
@@
When comparing the GPT models, all three models performed better with the WordLimit prompt, while yielding similar results with the iterative prompt (HGRWL). In terms of average length, the WordLimit prompt consistently gave the shortest summaries for all GPT models. GPT-4 and GPT-4o followed better the WordLimit prompt in obtaining a summary of less than 25 words than ChatGPT, which consistently exceeded the limit.
@@
Using DeepSeek-R1999https://api-docs.deepseek.com/guides/reasoning_model as our evaluation backbone, we evaluated the generated summaries against the human summarization guideline criteria (see Appendix B for details and prompts). We evaluated a subset of 20 samples from the DECODA dataset, with results shown in Table 4. To investigate the impact of length on quality, we include an additional column reporting summary length in our analysis.
However, we observed that even when the prompt explicitly provided guidelines for judging summaries and specified that only a score should be returned, many evaluation results contained not only scores but also explanatory justifications (sometimes in French and sometimes in English).
The GPT models’ generated summaries achieved strong scores across all four criteria. The best results for Faithfulness, Main Issues, and Resolution came from 4-HGRWL, 3.5-WL and 3.5-HGRWL respectively, while the Reference performed best on Sub-Issues. Considering the average scores across all criteria, 4-HGRWL achieved the highest values, while BARThez achieved the lowest.
| Faith. (VN) | Main Iss. (VN) | Sub-Iss. (VN) | Resol. (VN) | Avg. (VN) | Length | |
| Reference | 3.651.09 (20) | 4.351.09 (20) | 3.851.27 (20) | 3.151.81 (20) | 3.751.32 (20) | 24.8517.19 |
| 3.5-WL | 4.000.86 (20) | 4.600.75 (20) | 3.601.10 (20) | 3.351.73 (20) | 3.891.11 (20) | 44.0023.86 |
| 3.5-HGRWL | 3.201.47 (20) | 4.001.30 (20) | 3.751.21 (20) | 3.951.15 (20) | 3.731.28 (20) | 38.4010.18 |
| 4-WL | 4.001.12 (20) | 4.400.82 (20) | 3.421.39 (19)† | 2.451.70 (20) | 3.571.26 (20) | 23.35 3.08 |
| 4-HGRWL | 4.150.93 (20) | 4.450.76 (20) | 3.800.83 (20) | 3.501.54 (20) | 3.981.02 (20) | 25.20 3.30 |
| BARThez | 1.701.13 (20) | 2.001.65 (20) | 1.420.84 (19)‡ | 1.601.27 (20) | 1.681.22 (20) | 18.05 5.04 |
@@ 4.3 Manual Error Analysis
We examined the data points where GPT-generated summaries exhibit the greatest discrepancy, characterized by low ROUGE scores but high BERTScore values. For comparison, we selected predictions from two models: 4-WL (simple WordLimit prompt) and 4-HGRWL, which incorporates human summarization guidelines as an intermediate step. For DialogSum, we show top discrepancies from 4-WL (Table 6) and 4-HGRWL (Table 15). For DECODA, we present 4-WL results (Table 7) and 4-HGRWL (Table 16), both translated from French.
@@
We first analyze the summaries generated by 4-WL. In DialogSum examples with the greatest discrepancies, the references preserve specific mentions such as #Person1# and #Person2#, retaining named entities while summarizing the conversation. In contrast, GPT-4 predictions often provide a higher-level generalization, using generic terms like "two friends" or "employees" instead of specific tags. This indicates that GPT-generated summaries do not meet the following guideline: "3) Preserve important named entities within the conversation" when using a simplified prompt.
In addition, GPT-4 predictions do not match the references concerning the guideline emphasizing "Intent Identification".101010”Rather than merely summarizing the consequences of dialogues, you should also describe speakers intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation.” For instance, in the second example, the prediction references only Person 2, omitting the intent of Person 1, who is interviewing the other individual. In the fourth example, neither Person 1 nor Person 2 is mentioned, and their intents are absent, as the prediction reduces the dialogue to a general remark about the weather. In the fifth example, the prediction fails to follow the "use person tags" rule, summarizing the conversation by identifying the interlocutors as "A Mexican businesswoman" and "her American-Colombian taxi driver". These variations probably influenced the quantitative results, as evidenced by the disparity between high BERTScore values, which indicate strong semantic similarity, and low ROUGE-L scores, which reflect minimal lexical overlap. This contrast highlights the tendency of GPT-generated summaries using a simple length-constrained prompt to capture the essential meaning and key information of the conversation while diverging from the reference summaries in their specific wording and structural composition.
For the DECODA dataset, the discrepancy may stem from the distinct characteristics of the reference synopses. These summaries focus strictly on issues and resolutions, whereas GPT-generated predictions often adopt a more individual-centered perspective, using terms such as "un organisateur (an organizer)", "une cliente (a client)", and "un représentant (a representative)", etc. This difference results in relatively high BERTScore values, reflecting strong semantic similarity, but low ROUGE-L scores due to minimal word overlap with the references. Additionally, reference synopses are often extremely brief and sometimes incomplete sentences, as seen in the fourth example: "circulation du T2" (T2 circulation). Furthermore, all samples except the fourth fail to conclude the resolution in the generated summaries.
@@
Next, we examined the generated summaries of 4-HGRWL. On the DialogSum dataset, we observe better retention of named entities and personal tags (e.g., Person 1) compared to 4-WL, indicating that the intermediate HGR step helps the model to follow specific rules from the human summarization guidelines. However, the issue of overlooking "Intent Identification" persists, as prediction often focus on one individual’s intent while omitting the other’s, whereas references typically indicate both interlocutors’ intents.
For DECODA, similar to the 4-WL output, most generated summaries employ an individual-centered perspective that creates divergence in -gram overlap and semantic similarity. Although reference summaries maintain conciseness, model predictions tend toward detail while frequently omitting critical "Resolution" elements. For example, the first sample omits "Transfer to the relevant department" despite this being an explicit requirement in the summarization guidelines.
In summary, comparing cases with the greatest discrepancy (low ROUGE-L but high BERTScore) between 4-WL and 4-HGRWL, the latter adheres better to guidelines for named entities and personal tags. However, both models share similar shortcomings: in DialogSum, they overlook "Intent Identification"; in DECODA, they use an individual-centered perspective and omit "Resolution".